OpenStudy (anonymous):

98x + 133y = 427 Specify particular the answers (x,y) where x>0 and y>0 I get the answer (-244,183) or the general answer x= 19n-244 y= 183-14n I don't really get the x>0, y>0 part, could anyone guide me through it?

OpenStudy (anonymous):

they want u to find n's for which x and y are positive so u need to solve this inequalities \[19n-244>0\]\[183-14n>0\]

OpenStudy (anonymous):

So the answer should just be where \[n >\frac{ 244 }{ 19 }\] or \[n <\frac{ 183 }{ 14 }\] Right?

OpenStudy (anonymous):

not or ... it is and...

OpenStudy (anonymous):

thats right ...just note that they are simultaneous

OpenStudy (anonymous):

So the answer lies within the interval \[\frac{ 244 }{ 19}< n < \frac{ 183 }{14}\] ?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

just one number satisfies that inequalities :)

OpenStudy (anonymous):

13, thanks alot for your help! :D

OpenStudy (anonymous):

no problem